A contractor estimated that 30 ft^3 of cement is required to make a concrete floor that measure 90 ft^2. Using this estimate, how much cement is required to make a concrete floor that measures 120ft^2?
Answer by checkley77(12844)

The answer of 10ft^2 doesn't make sense.
You are asked to increase the area by 33%.
Therefore you'll nned 33% more cement ASSUMING the depth of the cement is the same.
Thus the amount of cement will be 33% greater or:
90=x/120
90x=30*120
90x=3,600
x=3,600/90
x=40 ft^3 of cement will be required.
You asked how much cement is requird. The answer is 30Yds^2
IF you had asked how much MORE cement would be required then 10^2 would be correct.